OpenStudy (anonymous):

cosΘ + cos ( Θ+ 2π) + cos ( Θ+ 4π), cos(Θ) = -0.6

mathslover (mathslover):

cos(a+b) = ?

mathslover (mathslover):

cos(A + B) = cosAcosB - sinAsinB M i correct @kikojuarez12 ?

mathslover (mathslover):

@kikojuarez12 please reply

OpenStudy (anonymous):

find the value of cosΘ + cos ( Θ+ 2π) + cos ( Θ+ 4π), cos(Θ) = -0.6

OpenStudy (anonymous):

fin the value of this cosΘ + cos ( Θ+ 2π) + cos ( Θ+ 4π)

OpenStudy (anonymous):

cos ( Θ+ 2π) assuming that is theta cos ( Θ+ 2π) =cos ( Θ) 2π is the same as 360 degrees thus Θ+ 2π= Θ

mathslover (mathslover):

2pi = 0 pi = 0 ? @completeidiot is that possible ?

OpenStudy (anonymous):

sigh... @mathslover you need to remember these are angles